OpenStudy (anonymous):

HELP how do I solve this in proof form? (1-sinx/cos)+(cosx/1/sinx)

OpenStudy (anonymous):

Is your question to solve this? \[1-\frac{\sin{x}}{\cos{x}}+\frac{\cos{x}}{\frac{1}{\sin{x}}} = 0\]

OpenStudy (anonymous):

no its 1-sinx/cos)+(cosx/1/sinx) =2secx

OpenStudy (anonymous):

Can you write it out using the equation editor for clarity please?

OpenStudy (anonymous):

\[\frac{ 1-\sin x}{ cosx }+ \frac{ cosx }{ 1-sinx }= 2 secx\]

OpenStudy (anonymous):

First things first, can you see how rewriting secx might simplify the problem?

OpenStudy (anonymous):

it want's to know how to reach 2secx from the given equation

OpenStudy (mathmale):

Perhaps. But almost all of this equation is in terms of sine and cosine. Wouldn't it be helpful to convert that sec x to an equivalent form in terms of cos x?

myininaya (myininaya):

Compare the sides. You need to get from two terms to one term. To do that combine the fractions by finding a common denominator